OpenStudy (anonymous):

Help Metal!!! Find the unknown side length for the rectangle. Area = 1.5 m2

OpenStudy (mathstudent55):

Is there a figure? BTW, metal is a substance such as gold, iron, aluminum, etc. Medal is an award.

OpenStudy (anonymous):

yes it is a square

OpenStudy (anonymous):

how do u metal

OpenStudy (mathstudent55):

Click "Best Response" in a response of the person you want to give the medal to.

OpenStudy (mathstudent55):

If the figure is a square, then all four sides have the same length, L. \(A = LW = LL = L^2\) \(L^2 = A\) \(L^2 = 1.5~m^2\) \(L = \sqrt{1.5~m^2}\) \(L = \sqrt{1.5} ~m\) The answer above is exact. Below is an approximation. \(L = \sqrt{1.5}~m \approx 1.225~ m\) (The symbol \(\approx\) is read "is approximately equal to.")